OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of s to the eighth power + the square root of 25 times s to the eighth power + 2 the square root of s to the eighth power − the square root of s to the fourth power 8s4 28 times s to the fourth power times the square root of s 8s4 − s2 28 times s to the fourth power times the square root of s − s2

OpenStudy (anonymous):

\[\sqrt{s ^{8}} + \sqrt{25s ^{8}} + 2\sqrt{s ^{8}} - \sqrt{s ^{4}}\]

OpenStudy (jhannybean):

\[\large \sqrt{s ^{8}} + \sqrt{25s ^{8}} + 2\sqrt{s ^{8}} - \sqrt{s ^{4}}\]You can rewrite this equation. You should already know that a square root, \(\sqrt{x}\) can be rewritten as \(x^{1/2}\) Taking that property, lets write all our terms to the half power. This will make it easier for us to evaluate.

OpenStudy (jhannybean):

\begin{align} \large \sqrt{s ^{8}} + \sqrt{25s ^{8}} + 2\sqrt{s ^{8}} - \sqrt{s ^{4}} &= \large s^{8/2} +(25^{1/2})(s^{8/2}) +2(s^{8/2})-s^{4/2} \\ &= \large s^4 +(5)(s^4)+2(s^4)-s^2 \\ &=\large s^4 +5s^4 +2s^4 -s^2 \\ &=\large 6s^4 +2s^4 -s^2 \\ \end{align}

OpenStudy (anonymous):

Thank you (: But that answer isnt in my multiple choice answers?

OpenStudy (jhannybean):

What do you get when you simplify this?

OpenStudy (anonymous):

8s^4 -s^2

OpenStudy (jhannybean):

There you go :)

OpenStudy (anonymous):

Yay! Thanks! Can you help with more?

OpenStudy (jhannybean):

I can try.